OpenStudy (anonymous):

Are these series absolutely convergent, conditionally convergent, or divergent?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ (-1)^{n-1} }{ 4n-5 }\]

OpenStudy (anonymous):

I think we could use the alternating series test for this one.

ganeshie8 (ganeshie8):

Yes that will work!

OpenStudy (anonymous):

Okay awesome!

OpenStudy (anonymous):

I think it might be conditionally convergent... since b_1 = -1, b_2 = 1/3, and b_3 = 1/7

OpenStudy (anonymous):

b_n+1 <= b_n is not fully satisfied.

OpenStudy (anonymous):

Would you agree?

ganeshie8 (ganeshie8):

it is indeed conditionally convergent, but your reasoning is wrong. "b_n+1 <= b_n is not fully satisfied" has nothing to do with conditional convergence

OpenStudy (anonymous):

Oh... oops

ganeshie8 (ganeshie8):

Could you show your complete work if psble

OpenStudy (anonymous):

The Alternating Series Test shows that a series must satisfy (1) b_n+1 >= b_n for all n (2) lim as n approaches infinity =0. b_n = 1/ (4n -5) in this case.

OpenStudy (anonymous):

b_1 = -1, b_2 = 1/3, b_3= 1/7. Thus the series does not completely satisfy b_n+1 >= b_n

OpenStudy (anonymous):

However it does satisfy the second part of the Alternating Series Test... lim as n approaches infinity of 1/4n -5 is zero.

OpenStudy (anonymous):

We note that the series is convergent. Now we can go to another test to see if it is absolutely convergent.

OpenStudy (anonymous):

Oh whoops, wrong notation. Sorry

ganeshie8 (ganeshie8):

I thought alternating series test requires the sequence to be "decreasing" : b_n+1 \(\color{red}{\lt}\) b_n right ?

OpenStudy (anonymous):

b_ n+1 <= b_n

OpenStudy (anonymous):

Well doesn't |b_n| converge? So wouldn't it be absolutely convergent?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

No, maybe lets not talk about absolute convergence yet. If we forget about the first term, the rest of the sequence is strictly decreasing, yes ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

So does the series for n > 1 meet all the requirements of alternating series test ?

OpenStudy (anonymous):

Nope!

ganeshie8 (ganeshie8):

why nope ?

ganeshie8 (ganeshie8):

just forget about the first term and look at the remaining series

OpenStudy (anonymous):

OH sorry, I thought you wrote n>=1. Yes

ganeshie8 (ganeshie8):

\[\sum_{n=2}^{\infty} \frac{ (-1)^{n-1} }{ 4n-5 }\] So by alternating series test, this series converges.

OpenStudy (anonymous):

Yep! Sounds good so far

ganeshie8 (ganeshie8):

Next see if the absolute value of the series also converges : \[\sum_{n=2}^{\infty} \left|\frac{ (-1)^{n-1} }{ 4n-5 }\right|\]

ganeshie8 (ganeshie8):

thats same as \[\sum_{n=2}^{\infty} \frac{ 1 }{ 4n-5 }\]

ganeshie8 (ganeshie8):

\[\frac{ 1 }{ 4n-5 }\gt \dfrac{1}{4n}\] Since the harmonic series \(\sum\limits_{n\ge 1}\frac{1}{4n}\) diverges, the absolute value series \(\sum\limits_{n\ge 2}\frac{1}{4n-5}\) also diverges by comparison test.

OpenStudy (loser66):

.

ganeshie8 (ganeshie8):

We therefore the say the series \(\sum_{n=1}^{\infty} \frac{ (-1)^{n-1} }{ 4n-5 }\) is "conditionally convergent".

ganeshie8 (ganeshie8):

we say it is conditionally convergent because of below two things : 1) the actual series converges 2) the absolute value of series diverges

OpenStudy (anonymous):

but the actual series does not converge... when n=1... it does not satisfy the alternating series test

ganeshie8 (ganeshie8):

forget about first term, it doesn't affect the convergence/divergence of series

OpenStudy (anonymous):

oh since the first term is such a small term that it will not greatly affect the convergence and divergence of the series?

OpenStudy (anonymous):

not a small term but just one term that doesn't satisfy the alternating series test.

ganeshie8 (ganeshie8):

you can forget about first "any finite" number of terms, they partial sum evaluates to some finite number, so they don't affect the convergence/divergence of actual series.

ganeshie8 (ganeshie8):

for example, you can forget about first 100 terms you could also forget first 10^10 terms

ganeshie8 (ganeshie8):

because the sum evaluates to some "finite number", however big it is..

OpenStudy (anonymous):

gotcha.

OpenStudy (anonymous):

Okay I think I'm getting the hang of this... could you please check my work for the next two series?

ganeshie8 (ganeshie8):

but the word "forget first few terms" doesn't look mathematical im just using it to make it easy for you to see what i mean..

OpenStudy (anonymous):

Yes, I see

ganeshie8 (ganeshie8):

In your homework, justify your reasoning like this :

OpenStudy (anonymous):

We know that the first finite number of the terms in the series evaluates to some finite number. Thus this does not affect the convergence or divergence of the series. We can now evaluate starting from n=2.

ganeshie8 (ganeshie8):

\[\sum_{n=1}^{\infty} \frac{ (-1)^{n-1} }{ 4n-5 }=-1+\color{blue}{\sum_{n=2}^{\infty} \frac{ (-1)^{n-1} }{ 4n-5 }}\] Since the blue series converges by alternating series test, adding a constant "-1" to the series still gives a converging series. Therefore the overall series on left hand side also converges.

ganeshie8 (ganeshie8):

"We know that the first finite number of the terms in the series evaluates to some finite number" this looks good !

OpenStudy (anonymous):

OOHH! Wow, I remember that

OpenStudy (anonymous):

Okay here is the next series...

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ 5n -1 }{ 3n +2 }\]

OpenStudy (anonymous):

If you can help :)

ganeshie8 (ganeshie8):

This one is easy try limit test

OpenStudy (anonymous):

approaches 1

ganeshie8 (ganeshie8):

nope

OpenStudy (anonymous):

5/3

OpenStudy (anonymous):

I think I'm overthinking this xD

ganeshie8 (ganeshie8):

the limit of the sequence is 5/3, which is not 0. so the series doesn't converge by limit test.

OpenStudy (anonymous):

Would you account for the (-1)^n part when taking the limit test?

ganeshie8 (ganeshie8):

i don't see (-1)^n in your series ?

OpenStudy (anonymous):

Oh sorry, I forgot it in there

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} (-1)^{n} \frac{ 5n -1 }{ 3n +2 }\]

OpenStudy (anonymous):

Could we use the alternating series test for this? and how?

ganeshie8 (ganeshie8):

when limit does the job for you, don't go for alternating series test.

ganeshie8 (ganeshie8):

and yes we should work the limit \(\lim\limits_{n\to\infty}(-1)^n\frac{5n-1}{3n+2}\)

ganeshie8 (ganeshie8):

simply working \(\lim\limits_{n\to\infty}\frac{5n-1}{3n+2}\) wont do

OpenStudy (anonymous):

Yes agreed.

ganeshie8 (ganeshie8):

However that limit with (-1)^n is a bit tricky because it toggles between -1 and 1

OpenStudy (anonymous):

Maybe using the ratio test then? I'm having difficulty choosing the right method

ganeshie8 (ganeshie8):

No, limit test will do.

ganeshie8 (ganeshie8):

Just put ur reasoning nicely

ganeshie8 (ganeshie8):

\[\lim\limits_{n\to\infty}(-1)^n\frac{5n-1}{3n+2}\] Since \(\lim\limits_{n\to\infty}\frac{5n-1}{3n+2}\) evaluates to \(\frac{5}{3}\), it must be the case that \(\lim\limits_{n\to\infty}(-1)^n\frac{5n-1}{3n+2}\) toggles between \(\frac{5}{3}\) and \(-\frac{5}{3}\), which is NOT 0. Therefore the corresponding series \(\sum_{n=1}^{\infty} (-1)^{n} \frac{ 5n -1 }{ 3n +2 }\) doesn't converge by limit test.

ganeshie8 (ganeshie8):

That will do

OpenStudy (anonymous):

I know this is wrong but why would the limit not be zero? would 5/3 and -5/3 cancel each other out?

ganeshie8 (ganeshie8):

You're confusing "limit of a sequence" with "series"

OpenStudy (anonymous):

Oh I see now.. oops

ganeshie8 (ganeshie8):

\[\lim\limits_{n\to\infty}(-1)^n\frac{5n-1}{3n+2}\] For large values of \(n\), the sequence \(a_n = (-1)^n\frac{5n-1}{3n+2}\) spits out terms that are very close to either \(\frac{5}{3}\) or \(-\frac{5}{3}\), yes ?

OpenStudy (anonymous):

Yes.. but the limit is oscillating so it is divergent.

OpenStudy (anonymous):

Okay I have one more series that I need help on... :)

ganeshie8 (ganeshie8):

Wil try, ask..

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ \sqrt{n-1} }{ n^{2} -6 } * (-1)^{n+1}\]

OpenStudy (anonymous):

Could we just use the limit test like last time? So it is divergent then...

ganeshie8 (ganeshie8):

how do you know it diverges ?

OpenStudy (anonymous):

Oh no, we could use the ratio test.. so that the limit of sqrt{n-1} / n^2 -6 is 0... thus it absolutely converges.

ganeshie8 (ganeshie8):

Yes this converges absolutely!

ganeshie8 (ganeshie8):

but i think your professor wants you to use alternating series test and show that the actual series converges first. after that you can use ratio test for testing convergence of absolute value sereis

OpenStudy (anonymous):

Alrighty! I'll make sure to do that then. Thank you ~!

ganeshie8 (ganeshie8):

yw:)